# A Vision for Trustworthy, Fair, and Efficient Socio-Technical Control using Karma Economies

### v1

@@ 2.4 Pillar 4: A future built on Nash welfare
For fairness and efficiency, which we discuss jointly under the more general term of social welfare, one should be mindful of the relevant information the urgency processes are meant to portray. If there is no reason to grant time-persistent priorities to some users over others, the urgency scale should once again play no role. If instead the urgency processes are meant to describe dynamics irrespective of scale, there is room to unequivocally promote both fairness and efficiency. Figure 5(b) shows an example of two urgency processes with the ‘same scale’, in the sense that they have the same long-run average urgency, but different dynamic natures. This could represent two types of users: one that is frequently late and in a hurry (Figure 5(b)-left) and another that is less often late (Figure 5(b)-right). It would be both fair and efficient to prioritize a user of the second type over the first in the few times it has high urgency: the second user requests priority less frequently; and in turn, the urgency of the second user relative to other days is higher than the urgency of the first relative to other days.
The above discussion can be summarized in requiring that urgency processes be interpersonally non-comparable. Fortunately, there is a social welfare function that elegantly encodes the possible efficiency gains under this assumption, that is, the Nash welfare function, given by {mdframed}[backgroundcolor=red!10]
$\textup{NW}\left(\bm{\chi}\right)=\prod_{i\in{\mathcal{N}}}\left(\bar{r}_{i}\left(\bm{\chi}\right)-\bar{r}_{i}\left(\bm{\chi^{0}}\right)\right)^{w_{i}}.$ (NW)
@@
In (NW), $\bm{\chi}\in\Delta({\mathcal{X}}^{T})$ denotes a probabilistic allocation over all times $t\in{\mathcal{T}}$, $\bar{r}_{i}\left(\bm{\chi}\right)$ is the expected average reward of user $i\in{\mathcal{N}}$ under $\bm{\chi}$, and $\bm{\chi^{0}}\in\Delta({\mathcal{X}}^{T})$ is a benchmark or status-quo allocation in which all users are worst-off. The importance of this benchmark is discussed in what follows. In the running example of allocating priority lanes, the benchmark could be that no control measure is introduced, i.e., all lanes are general purpose. Finally, $\bm{w}=\left(w_{i}\right)_{i\in{\mathcal{N}}}\in{\mathbb{R}}_{>0}^{n}$ denote exogenous user-specific priority weights or access rights. Most commonly, $w_{i}=1$ for all $i\in{\mathcal{N}}$ means that all users have equal rights; but there are cases in which unequal rights could be warranted, e.g., prioritizing emergency care givers over regular citizens in accessing fast lanes.
@@
Nash welfare is a classical social welfare function that has been portrayed to ‘strike a good trade-off between fairness and efficiency’ (Caragiannis et al., 2019; Bertsimas et al., 2012). This motivation can be misleading because ‘trading off’ fairness and efficiency implicitly assumes interpersonal comparability. In what follows, we argue that Nash welfare jointly encodes fairness and efficiency, rather than trades off between the two.
{mdframed}
@@
Parallel to studying the possibilities and impossibilities of designing allocation rules ala Theorems 12, social choice theory investigates how to meaningfully aggregate individual preferences into social welfare functions (Roberts, 1980). Examples of social welfare functions are the utilitarian sum of rewards, the egalitarian minimum reward, and Nash welfare. Shilov et al. (2025) provides a comprehensive overview of results in this space tailored to a control audience, and we provide an informal summary of the basic elements here. The choice of social welfare function is guided by the following set of basic properties or axioms that the function should satisfy:
  • Pareto principle (P): If allocation $x$ Pareto dominates allocation $y$, then $x$ should attain higher social welfare than $y$;
  • Independence of irrelevant alternatives (IIA): The social ordering of two allocations $x$ and $y$, i.e., which allocation of the two attains higher social welfare, depends only on the users’ rewards at $x$ and $y$ and not on the rewards of other allocations777(IIA) is the most often critiqued axiom of social choice theory. However, in addition to its basic premise, (IIA) has an important connection to strategy-proofness: under additional mild conditions, an allocation rule is strategy-proof if and only if it maximizes a social welfare function satisfying (IIA) (Satterthwaite, 1975)..
  • Comparability: Assumptions on how to compare the rewards of different users range from the weakest requirement of ordinal non-comparability (ONC), which considers rewards as ordinal and not interpersonally comparable; to the strongest requirement of perfect comparability (PC), which considers that rewards are perfectly comparable on a common absolute scale.
Depending on the adopted comparability requirement, there is either no possibility, a unique possibility, or a restricted set of possibilities to jointly satisfy (P) and (IIA) (Shilov et al., 2025, Theorem 1). Remarkably, the weakest comparability requirement for which it becomes possible to satisfy (P) and (IIA) is cardinal non-comparability (CNC), which postulates that rewards are cardinally intrapersonally comparable but interpersonally non-comparable. Under (CNC), the unique social welfare function satisfying (P) and (IIA) is the Nash welfare function (NW). The worst-case benchmark allocation $\bm{x^{0}}$ is important for this result, namely, without such a benchmark it is impossible to satisfy (P), (IIA), and (CNC). Intuitively, the benchmark allows making meaningful interpersonal trade-offs without making interpersonal comparisons; by comparing intrapersonal relative improvements to the benchmark. In light of the preceding discussion to regard heterogeneous urgency processes as interpesonally non-comparable, social choice theory thus provides strong foundations for adopting (NW) as the social welfare function in dynamic resource allocation problems.
{mdframed}
@@
A parallel and independent motivation for Nash welfare stems from bargaining theory. In fact, the name Nash welfare attributes its origins to the Nash bargaining solution (Nash et al., 1950) [^8] The Nash bargaining solution is not to be confused by the famous Nash equilibrium, and in fact Nash welfare has little to nothing to do with the Nash equilibrium.. In the bargaining problem, two or more users must come to an agreement on an allocation $x$, and if negotiations fail, they must fall back on the status quo or default option $x^{0}$. The bargaining problem is ill-defined because there are often many allocations that would benefit all users, but could differ significantly in the distribution of gains. The Nash bargaining solution, which maximizes Nash welfare, is widely regarded as the fair outcome of the bargain: in the case of dividing monetary gains, it leads to splitting the surplus equally after compensating the users for the opportunity cost of forfeiting the default option $x^{0}$. Moreover, if the bargaining is performed sequentially or repeatedly, it has been shown that the Nash bargaining solution emerges as the unique equilibrium (Rubinstein, 1982; Binmore, 2005). Based on this principle, it has been suggested that moral principles are in fact social contracts that have evolved to maximize Nash welfare (André et al., 2022).
One could adopt a bargaining interpretation of socio-technical control problems. Society must come to an agreement on what control scheme to adopt. In the case of NYC congestion pricing, c.f. Section 2.2, the negotiations lasted 73 years; perhaps they would have concluded quicker if one targeted maximum Nash welfare.
@@ 3 Dynamic Resource Allocation under Maximum Long-Run Nash Welfare (MLNW)
This is the first of two sections that introduce the machinery for operationalizing our vision for trustworthy, fair, and efficient socio-techical control. We formally introduce dynamic resource allocation problems (Section 3.1); before taking a centralized view on Maximum Long-Run Nash Welfare (MLNW) solutions to highlight their fairness and efficiency properties (Section 3.2). Then, in the subsequent Section 4, we show that MLNW can be attained in a decentralized and trustworthy manner using karma economies.
@@ 3.1 Problem formulation
We consider the problem of repeatedly allocating resources =${\mathcal{M}}=\{1,\dots,m\}$ (indexed $j$) of capacities $\bm{c}=\left(c_{j}\right)_{j\in{\mathcal{M}}}\in{\mathbb{N}}_{>0}^{m}$ to users ${\mathcal{N}}=\{1,\dots,n\}$ over an infinite time. The users have exogenous access rights $\bm{w}=(w_{i})_{i\in{\mathcal{N}}}\in{\mathbb{R}}_{>0}^{n}$ representing long-run priorities in accessing resources. The allocation at time-step $t\in{\mathbb{N}}$ is denoted by $\bm{x_{t}}\in{\mathcal{X}}=\left\{\bm{x}\in\{0,1\}^{n\times m};\;\sum_{i\in{\mathcal{N}}}x_{i,j}\leq c_{j},\;\forall j\in{\mathcal{M}}\right\}$, where $x_{t,i,j}=1$ (respectively, $x_{t,i,j}=0$) denotes that user $i$ is allocated (respectively, not allocated) resource $j$. The sub-allocation to user $i$ is denoted by $\bm{x_{t,i}}=\left(x_{t,i,j}\right)_{j\in{\mathcal{M}}}\in\{0,1\}^{m}$. In some settings, the resources are mutually exclusive, i.e., each user can be allocated at most one resource at a time, for which it must also hold that $\sum_{j\in{\mathcal{M}}}x_{t,i,j}\leq 1$ for all $i\in{\mathcal{N}}$.
@@
The urgency $u_{t,i}$ of each user follows a private, exogenous stochastic process that is independent across users and time, denoted by $\bm{\sigma_{i}}\in\Delta({\mathcal{U}}_{i})$ and satisfying $\sigma_{i}(u_{i})>0$ for all $u_{i}\in{\mathcal{U}}_{i}$. The dynamic resource allocation problem is thus specified by the tuple ${\mathfrak{D}}=\left\{{\mathcal{N}},{\mathcal{M}},\bm{w},\bm{c},\bm{r},\bm{r^{0}},\bm{{\mathcal{U}}},\bm{\sigma}\right\}$, where $\bm{r}=\left(r_{i,j}\right)_{i\in{\mathcal{N}},j\in{\mathcal{M}}}$, $\bm{r^{0}}=\left(r^{0}_{i,j}\right)_{i\in{\mathcal{N}},j\in{\mathcal{M}}}$, $\bm{{\mathcal{U}}}=\left({\mathcal{U}}_{i}\right)_{i\in{\mathcal{N}}}$, and $\bm{\sigma}=\left(\bm{\sigma_{i}}\right)_{i\in{\mathcal{N}}}$. We illustrate this notation on the following extension of our running example that incorporates the allocation of highway priority lanes at different times of the day.
{mdframed}
[backgroundcolor=green!10]
@@
[…]
@@ 3.2 The fairness and efficiency of Maximum Long-run Nash Welfare (MLNW)
{mdframed}
[backgroundcolor=red!10] MLNW Problem: $\max_{\bm{\chi}}$ $\sum_{i\in{\mathcal{N}}}w_{i}\log\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)$ (MLNW) subject to $\sum_{i\in{\mathcal{N}}}\sum_{u_{i}\in{\mathcal{U}}_{i}}\sigma_{i}(u_{i})\,\chi_{i,j}(u_{i})\leq c_{j},\;\forall j\in{\mathcal{M}},$ $0\leq\chi_{i,j}(u_{i})\leq 1,\;\forall i\in{\mathcal{N}},j\in{\mathcal{M}},u_{i}\in{\mathcal{U}}_{i}.$
@@
Let Assumption 1 hold. Then MLNW solutions $(\bm{\chi^{*}},\bm{\lambda^{*}},\bm{\eta^{*}})$ satisfy, for all $i\in{\mathcal{N}}$, $j\in{\mathcal{M}}$, $u_{i}\in{\mathcal{U}}_{i}$:
  1. 3.1
    $\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{*}>0$ and is unique;
  2. 3.2
[…]
  3. 3.3
    $\chi^{*}_{i,j}(u_{i})>0\Rightarrow\chi^{*}_{i,j}(u^{\prime}_{i})=1$, for all $u^{\prime}_{i}>u_{i}$;
  4. 3.4
    $\lambda^{*}_{j}>0$, and all resources are allocated at capacity;
  5. 3.5
    $u_{i}\left(r_{i,j}-r^{0}_{i,j}\right)\leq\cfrac{\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{*}}{w_{i}}\left(\lambda^{*}_{j}+\cfrac{\eta^{*}_{i,j}(u_{i})}{\sigma_{i}(u_{i})}\right)$;
  6. 3.6
    $\chi^{*}_{i,j}(u_{i})>0\Rightarrow u_{i}\left(r_{i,j}-r^{0}_{i,j}\right)=\cfrac{\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{*}}{w_{i}}\left(\lambda^{*}_{j}+\cfrac{\eta^{*}_{i,j}(u_{i})}{\sigma_{i}(u_{i})}\right)$;
  7. 3.7
    $\sum\limits_{u_{i}\in{\mathcal{U}}_{i}}\sigma_{i}(u_{i})\sum\limits_{j\in{\mathcal{M}}}\chi^{*}_{i,j}(u_{i})\,\lambda^{*}_{j}+\sum\limits_{j\in{\mathcal{M}}}\sum\limits_{u_{i}\in{\mathcal{U}}_{i}}\eta^{*}_{i,j}(u_{i})=w_{i}$.
Moreover, if the resources are mutually exclusive, then Propositions 3.13.2 hold, Propositions 3.53.7 hold with non-resource specific multipliers $\eta^{*}_{i,j}(u_{i})=\eta^{*}_{i}(u_{i})$, and it holds for all $i\in{\mathcal{N}}$, $j\in{\mathcal{M}}$, $u_{i}\in{\mathcal{U}}_{i}$:
  1. 3.3
    $\chi^{*}_{i,j}(u_{i})>0\Rightarrow\sum_{j^{\prime}\in{\mathcal{M}}}\chi^{*}_{i,j^{\prime}}(u^{\prime}_{i})=1$, for all $u^{\prime}_{i}>u_{i}$;
  2. 3.4
    $\lambda^{*}_{j}=0\Rightarrow\sum_{j^{\prime}\in{\mathcal{M}}}\chi^{*}_{i,j^{\prime}}(u_{i})=1$, for all $i\in{\mathcal{N}}_{j}$, $u_{i}\in{\mathcal{U}}_{i}$.

### v2

@@ 2.4 Pillar 4: A future built on Nash welfare
For fairness and efficiency, which we discuss jointly under the more general term of social welfare, one should be mindful of the relevant information the urgency processes are meant to portray. If there is no reason to grant time-persistent priorities to some users over others, the urgency scale should once again play no role. Figure 5(b) shows an example of two urgency processes with the ‘same scale’, in the sense that they have the same long-run average urgency, but different dynamic natures. Despite having ‘same scale’, it is not straightforward to compare urgency levels of different processes, since these levels are experienced at different frequencies. Nonetheless, there is room to unequivocally promote both fairness and efficiency without making interpersonal comparisons of urgency. For example, it is natural to prioritize urgency $6$ in Figure 5(b)-right over urgency $3$ in Figure 5(b)-left, not because $6>3$ in an absolute sense, but because urgency $6$ occurs less frequently; and in turn, forms a more intense intrapersonal deviation relative to other days than urgency $3$ does.
@@
The above discussion can be summarized in requiring that urgency be interpersonally non-comparable. Fortunately, there is a social welfare function that elegantly encodes the possible efficiency gains under this assumption, that is, the Nash welfare function, given by $\textup{NW}\left(\bm{\chi}\right)=\prod_{i\in{\mathcal{N}}}\left(\bar{r}_{i}\left(\bm{\chi}\right)-\bar{r}_{i}\left(\bm{\chi^{0}}\right)\right)^{w_{i}}.$ (NW) In (NW), $\bm{\chi}\in\Delta({\mathcal{X}}^{T})$ denotes a probabilistic allocation over all times $t\in{\mathcal{T}}$, $\bar{r}_{i}\left(\bm{\chi}\right)$ is the expected average reward of user $i\in{\mathcal{N}}$ under $\bm{\chi}$, and $\bm{\chi^{0}}\in\Delta({\mathcal{X}}^{T})$ is a benchmark or status-quo allocation in which all users are worst-off. In the running example of allocating priority lanes, the benchmark could be that no control measure is introduced, i.e., all lanes are general purpose, cf. Example 4. Finally, $\bm{w}=\left(w_{i}\right)_{i\in{\mathcal{N}}}\in{\mathbb{R}}_{>0}^{n}$ denote exogenous user-specific priority weights or access rights. Most commonly, $w_{i}=1$ for all $i\in{\mathcal{N}}$ means that all users have equal rights; but there are cases in which unequal rights could be warranted, e.g., prioritizing emergency care givers over regular citizens in accessing fast lanes.
@@
Nash welfare is a classical social welfare function that has been portrayed to ‘strike a good trade-off between fairness and efficiency’ (Caragiannis et al., 2019; Bertsimas et al., 2012). This motivation can be misleading because ‘trading off’ fairness and efficiency implicitly assumes interpersonal comparability [^7] Namely, the sum of rewards that is most commonly used to define efficiency, as well as most existing fairness measures (e.g., minimum reward, Gini index, etc.), assume interpersonal comparability.. In what follows, we argue that Nash welfare jointly encodes fairness and efficiency, rather than trades off between the two.
@@
Parallel to studying the possibilities and impossibilities of designing allocation rules ala Theorems 12, social choice theory investigates how to meaningfully aggregate individual preferences into social welfare functions (Roberts, 1980). Examples of social welfare functions are the utilitarian sum of rewards, the egalitarian minimum reward, and Nash welfare. Shilov et al. (2025) provides a comprehensive overview of results in this space tailored to a control audience, and we provide an informal summary of the basic elements here. The choice of social welfare function is guided by the following set of basic properties or axioms that the function should satisfy: Pareto principle (P): If allocation $x$ Pareto dominates allocation $y$, then $x$ should attain higher social welfare than $y$; Independence of irrelevant alternatives (IIA): The social ordering of two allocations $x$ and $y$, i.e., which allocation of the two attains higher social welfare, depends only on the users’ rewards at $x$ and $y$ and not on the rewards of other allocations888(IIA) is the most often critiqued axiom of social choice theory. However, in addition to its basic premise, (IIA) has an important connection to strategy-proofness: under additional mild conditions, an allocation rule is strategy-proof if and only if it maximizes a social welfare function satisfying (IIA) (Satterthwaite, 1975).. Comparability: Assumptions on how to compare the rewards of different users range from the weakest requirement of ordinal non-comparability (ONC), which considers rewards as ordinal [^9] Ordinal rewards express rankings of allocations: they tell if an allocation is preferred over another, but not by how much. Cardinal rewards additionally express preference intensities in relative terms: for any three allocations $x$, $y$, $z$, they quantify the reward difference between $x$ and $y$ relative to the difference between $x$ and $z$; cf. Shilov et al. (2025) for the formal definitions. and not interpersonally comparable; to the strongest requirement of perfect comparability (PC), which considers that rewards are perfectly comparable on a common absolute scale. Depending on the adopted comparability requirement, there is either no possibility, a unique possibility, or a restricted set of possibilities to jointly satisfy (P) and (IIA) (Shilov et al., 2025, Theorem 1). Remarkably, the weakest comparability requirement for which it becomes possible to satisfy (P) and (IIA) is cardinal non-comparability (CNC), which postulates that rewards are cardinally intrapersonally comparable but interpersonally non-comparable. Under (CNC), the unique social welfare function satisfying (P) and (IIA) is the Nash welfare function (NW). The worst-case benchmark allocation $\bm{x^{0}}$ is important for this result, namely, without such a benchmark it is impossible to satisfy (P), (IIA), and (CNC). Intuitively, the benchmark circumvents this impossibility because it acts as a common point with respect to which relative intrapersonal improvements can be compared. In light of the preceding discussion to regard urgency of heterogeneous processes as interpesonally non-comparable, social choice theory thus provides strong foundations for adopting (NW) as the social welfare function in dynamic resource allocation problems.
@@
A parallel and independent motivation for Nash welfare stems from bargaining theory. In fact, the name Nash welfare attributes its origins to the Nash bargaining solution (Nash et al., 1950) [^10] The Nash bargaining solution is not to be confused by the famous Nash equilibrium, and in fact Nash welfare has little to nothing to do with the Nash equilibrium.. In the bargaining problem, two or more users must come to an agreement on an allocation $x$, and if negotiations fail, they must fall back on the status quo or default option $x^{0}$. The bargaining problem is ill-defined because there are often many allocations that would benefit all users, but could differ significantly in the distribution of gains. The Nash bargaining solution, which maximizes Nash welfare, is widely regarded as the fair outcome of the bargain: in the case of dividing monetary gains, it leads to splitting the surplus equally after compensating the users for the opportunity cost of forfeiting the default option $x^{0}$. Moreover, if the bargaining is performed sequentially or repeatedly, it has been shown that the Nash bargaining solution emerges as the unique equilibrium (Rubinstein, 1982; Binmore, 2005). Based on this principle, it has been suggested that moral principles are in fact social contracts that have evolved to maximize Nash welfare (André et al., 2022). One could adopt a bargaining interpretation of socio-technical control problems. Society must come to an agreement on what control scheme to adopt. In the case of NYC congestion pricing, c.f. Section 2.2, the negotiations lasted 73 years; perhaps they would have concluded quicker if one targeted maximum Nash welfare.
@@ 3 Dynamic Resource Allocation under Maximum Long-Run Nash Welfare (MLNW)
In this section we formally introduce dynamic resource allocation problems (Section 3.1) and then we take a centralized view on Maximum Long-Run Nash Welfare (MLNW) solutions to highlight their fairness and efficiency properties (Section 3.2).
@@ 3.1 Problem formulation
Our problem formulation adapts and extends canonical single-shot resource allocation models (Vazirani, 2007; Hylland and Zeckhauser, 1979) to consider infinite repetition and dynamic urgency. We consider the problem of repeatedly allocating resources ${\mathcal{M}}=\{1,\dots,m\}$ (indexed $j$) of capacities $\bm{c}=\left(c_{j}\right)_{j\in{\mathcal{M}}}\in{\mathbb{N}}_{>0}^{m}$ to users ${\mathcal{N}}=\{1,\dots,n\}$ over an infinite time. The users have exogenous access rights $\bm{w}=(w_{i})_{i\in{\mathcal{N}}}\in{\mathbb{R}}_{>0}^{n}$ representing long-run priorities in accessing resources. The allocation at time-step $t\in{\mathbb{N}}$ is denoted by $\bm{x_{t}}\in{\mathcal{X}}=\left\{\bm{x}\in\{0,1\}^{n\times m};\;\sum_{i\in{\mathcal{N}}}x_{i,j}\leq c_{j},\;\forall j\in{\mathcal{M}}\right\}$, where $x_{t,i,j}=1$ (respectively, $x_{t,i,j}=0$) denotes that user $i$ is allocated (respectively, not allocated) resource $j$. The sub-allocation to user $i$ is denoted by $\bm{x_{t,i}}=\left(x_{t,i,j}\right)_{j\in{\mathcal{M}}}\in\{0,1\}^{m}$. In some settings, the resources are mutually exclusive, i.e., each user can be allocated at most one resource at a time, for which it must also hold that $\sum_{j\in{\mathcal{M}}}x_{t,i,j}\leq 1$ for all $i\in{\mathcal{N}}$.
@@
The urgency $u_{t,i}$ of each user follows a private, exogenous stochastic process that is independent across users and time, denoted by $\bm{\sigma_{i}}\in\Delta({\mathcal{U}}_{i})$ and satisfying $\sigma_{i}(u_{i})>0$ for all $u_{i}\in{\mathcal{U}}_{i}$. The dynamic resource allocation problem is thus specified by the tuple ${\mathfrak{D}}=\left\{{\mathcal{N}},{\mathcal{M}},\bm{w},\bm{c},\bm{r},\bm{r^{0}},\bm{{\mathcal{U}}},\bm{\sigma}\right\}$, where $\bm{r}=\left(r_{i,j}\right)_{i\in{\mathcal{N}},j\in{\mathcal{M}}}$, $\bm{r^{0}}=\left(r^{0}_{i,j}\right)_{i\in{\mathcal{N}},j\in{\mathcal{M}}}$, $\bm{{\mathcal{U}}}=\left({\mathcal{U}}_{i}\right)_{i\in{\mathcal{N}}}$, and $\bm{\sigma}=\left(\bm{\sigma_{i}}\right)_{i\in{\mathcal{N}}}$. We illustrate this notation on the following extension of our running example.
@@
Consider now that the highway priority lanes are allocated repeatedly over the morning rush hour. Following Elokda et al. (2024c) that is based on the classical bottleneck model (Vickrey, 1969), suppose that every morning, $n=9000$ single occupancy vehicle users must traverse a highway of total capacity $60$ vehicles per minute; thus the morning rush hour lasts $150$ minutes. The highway is split into priority lanes of capacity $12$ vehicles per minute ($20\%$ of the total capacity), and general purpose lanes of capacity $48$ vehicles per minute. Let us discretize the departure times into $15$-minute intervals, such that the rush hour lasts over $10$ departure intervals. We are thus faced with the problem of repeatedly allocating ${\mathcal{M}}=\{1,\dots,10\}$ mutually exclusive resources, corresponding to the priority lanes at different departure intervals, whose capacities are $c_{j}=c=180$ for all $j\in{\mathcal{M}}$, that is, the free-flow capacity of the priority lanes over one interval. The departure intervals are not equally favorable to the users. Namely, all users $i\in{\mathcal{N}}$ wish to depart at interval $j^{*}=9$ (e.g., corresponding to 9:00am), and experience a reward of $r_{i,j}=r_{j}=-\max\{j^{*}-j,\,4\left(j-j^{*}\right)\}$, i.e., they incur a proportionally larger discomfort the further they depart from $j^{*}$, and are more sensitive towards late than early departures. In case a user is not allocated to the priority lanes at any interval, they must use the general purpose lanes, which gives a reward of $r^{0}_{i,j}=r^{0}=-8$ corresponding to the non-controlled equilibrium discomfort in those lanes [^11] The non-controlled equilibrium discomfort of the general purpose lanes additionally takes into account delays due to the build-up of queues. It coincides with the queue-free discomfort of departing in the earliest interval $j=1$, as well as the equilibrium discomfort if all lanes are not controlled, cf. Elokda et al. (2024c).. Notice that Assumption 1 holds in this setting with the exception of the earliest interval $j=1$ which gives the same reward as $r^{0}$. Moreover, suppose that all users have equal access rights, i.e., $w_{i}=1$ for all $i\in{\mathcal{N}}$.
@@ 3.2 The fairness and efficiency of Maximum Long-run Nash Welfare (MLNW)
MLNW Problem: $\max_{\bm{\chi}}$ $\sum_{i\in{\mathcal{N}}}w_{i}\log\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)$ (MLNW) subject to $\sum_{i\in{\mathcal{N}}}\sum_{u_{i}\in{\mathcal{U}}_{i}}\sigma_{i}(u_{i})\,\chi_{i,j}(u_{i})\leq c_{j},\;\forall j\in{\mathcal{M}},$ $0\leq\chi_{i,j}(u_{i})\leq 1,\;\forall i\in{\mathcal{N}},j\in{\mathcal{M}},u_{i}\in{\mathcal{U}}_{i}.$
@@
Let Assumption 1 hold. Then MLNW solutions $(\bm{\chi^{*}},\bm{\lambda^{*}},\bm{\eta^{*}})$ satisfy, for all $i\in{\mathcal{N}}$, $j\in{\mathcal{M}}$, $u_{i}\in{\mathcal{U}}_{i}$: 3.1 $\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{*}>0$ and is unique; 3.2 If two users $i$, $i^{\prime}$ are symmetric up to scales $\alpha_{i},\alpha_{i^{\prime}}>0$ [^14] Two users $i$, $i^{\prime}$ are called symmetric up to scales $\alpha_{i},\alpha_{i^{\prime}}>0$ if they share the same characteristics up to scaling of the urgency, i.e., $w_{i}=w_{i^{\prime}}$, $\bm{r_{i}}=\bm{r_{i^{\prime}}}$, $\bm{r^{0}_{i}}=\bm{r^{0}_{i^{\prime}}}$, $\bm{\phi_{i}}=\bm{\phi_{i^{\prime}}}$, and there exists a common urgency set $\mathcal{U}$} such that ${\mathcal{U}}_{i}=\alpha_{i}\,{\mathcal{U}}$ and ${\mathcal{U}}_{i^{\prime}}=\alpha_{i^{\prime}}\,{\mathcal{U}}$., then $\cfrac{\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{*}}{\alpha_{i}}=\cfrac{\left(\bar{r}_{i^{\prime}}-\bar{r}^{0}_{i^{\prime}}\right)^{*}}{\alpha_{i^{\prime}}}$; 3.3 $\chi^{*}_{i,j}(u_{i})>0\Rightarrow\chi^{*}_{i,j}(u^{\prime}_{i})=1$, for all $u^{\prime}_{i}>u_{i}$; 3.4 $\lambda^{*}_{j}>0$, and all resources are allocated at capacity; 3.5 $u_{i}\left(r_{i,j}-r^{0}_{i,j}\right)\leq\cfrac{\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{*}}{w_{i}}\left(\lambda^{*}_{j}+\cfrac{\eta^{*}_{i,j}(u_{i})}{\sigma_{i}(u_{i})}\right)$; 3.6 $\chi^{*}_{i,j}(u_{i})>0\Rightarrow u_{i}\left(r_{i,j}-r^{0}_{i,j}\right)=\cfrac{\left(\bar{r}_{i}-\bar{r}^{0}_{i}\right)^{*}}{w_{i}}\left(\lambda^{*}_{j}+\cfrac{\eta^{*}_{i,j}(u_{i})}{\sigma_{i}(u_{i})}\right)$; 3.7 $\sum\limits_{u_{i}\in{\mathcal{U}}_{i}}\sigma_{i}(u_{i})\sum\limits_{j\in{\mathcal{M}}}\chi^{*}_{i,j}(u_{i})\,\lambda^{*}_{j}+\sum\limits_{j\in{\mathcal{M}}}\sum\limits_{u_{i}\in{\mathcal{U}}_{i}}\eta^{*}_{i,j}(u_{i})=w_{i}$. Moreover, if the resources are mutually exclusive, then Propositions 3.13.2 hold, Propositions 3.53.7 hold with non-resource specific multipliers $\eta^{*}_{i,j}(u_{i})=\eta^{*}_{i}(u_{i})$, and it holds for all $i\in{\mathcal{N}}$, $j\in{\mathcal{M}}$, $u_{i}\in{\mathcal{U}}_{i}$: 3.3 $\chi^{*}_{i,j}(u_{i})>0\Rightarrow\sum_{j^{\prime}\in{\mathcal{M}}}\chi^{*}_{i,j^{\prime}}(u^{\prime}_{i})=1$, for all $u^{\prime}_{i}>u_{i}$; 3.4 $\lambda^{*}_{j}=0\Rightarrow\sum_{j^{\prime}\in{\mathcal{M}}}\chi^{*}_{i,j^{\prime}}(u_{i})=1$, for all $i\in{\mathcal{N}}_{j}$, $u_{i}\in{\mathcal{U}}_{i}$.